OpenStudy (anonymous):

(1-tan([theta])^(2))/sec([theta]) =1, sec([theta]) + 2sin([theta]) help me verify it

OpenStudy (anonymous):

Something wrong there...?

OpenStudy (anonymous):

YEAHHH SEMANGAT

OpenStudy (anonymous):

no i have to show the steps tpo verify it to make both sides equal

OpenStudy (anonymous):

This part /sec([theta]) =1, sec([theta]) + doesn't make sense.

OpenStudy (anonymous):

the / is divided by

OpenStudy (anonymous):

=1, sec([theta]) Please use equation editor or attach a pic.

OpenStudy (anonymous):

\[1-\tan \Theta ^{2}\div \sec \theta = \sec \theta + 2\sin \theta\]

OpenStudy (anonymous):

1-tan^2 = 1-(sec^2-1) = sec^2 +2 So then multiply through by sec and the rhs comes out to the same.